A shop sells a package of holiday gift wrap for $25.50. It contains five small rolls and three large rolls. A package of birthday wrap with three small rolls and two large rolls sells for $16.25. The manager wants to sell a new alloccasion assortment. It will include two small rolls and two large rolls of gift wrap. The price of this package should be based on the prices of the small and large rolls in the other gift wrap packages. How much should the store charge for the new assortment?
Find the equation(s) and solve using substitution or elimination.
Not in a text book!

Answer by ankor@dixie-net.com(12706)

A shop sells a package of holiday gift wrap for $25.50. It contains five small rolls and three large rolls. A package of birthday wrap with three small rolls and two large rolls sells for $16.25. The manager wants to sell a new alloccasion assortment. It will include two small rolls and two large rolls of gift wrap. The price of this package should be based on the prices of the small and large rolls in the other gift wrap packages. How much should the store charge for the new assortment?
:
Find the equation(s) and solve using substitution or elimination.
:
Let L = price of large rolls, S = price of small rolls
:
Write an equation for each sentence:
:
"A shop sells a package of holiday gift wrap for $25.50. It contains five small rolls and three large rolls."
5S + 3L = 25.50
:
"A package of birthday wrap with three small rolls and two large rolls sells for $16.25."
3S + 2L = 16.25
:
Using the elimination method:
5s + 3L = 25.50 * 2 equals: 10s + 6L = 51.00
3s + 2L = 16.25 * 3 equals: 9s + 6L = 48.75
................................................----------------- Subtract
.....................................................1s + 0L = 2.25
..............................................................s = $2.25
:
Find the price of the large rolls (L)
3S + 2L = 16.25
3(2.25) + 2L = 16.25
6.75 + 2L = 16.25
2L = 16.25 - 6.75
L = 9.50/2
L = $4.75
:
Check our solutions: 5(2.25) + 3(4.75) = 25.50
:
"New package will include two small rolls and two large rolls of gift wrap."
:
New package cost: 2(2.25) + 2(4.75) = $14.00